OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + 2y = 6 2x + 3y = −6 6 −3 3 −6

OpenStudy (the_fizicx99):

Hmm, 3x + 2y = 6 2x + 3y = −6 I'm just going to use substitution as it seems easier. 3x + 2y = 6 -3x -3x 2y = -3x + 6 /2 /2 /2 y = -3/2x + 3 2x + 3(-1.5 + 3) = −6 2x -4.5 + 9 = -6 2x + 4.5 = -6 -4.5 -4.5 2x = -10.5 /2 /2 x = -5.25 Plug in -5.25 for x and solve for y

OpenStudy (anonymous):

-6?@tHe_FiZiCx99

OpenStudy (wolf1728):

A) 3x + 2y = 6 B) 2x + 3y = −6 Multiply Equation A by -1.5 A) -4.5x - 3y = -9 then add to B B) 2x + 3y = −6 -1.5x = -15 x = 6 A) 3•6 + 2y = 6 18 + 2y = 6 2y = -12 y = -6

OpenStudy (the_fizicx99):

Yes, -6

OpenStudy (anonymous):

Thanks @tHe_FiZiCx99 and @wolf1728 !!!

OpenStudy (wolf1728):

u r welcome :-)

OpenStudy (the_fizicx99):

I see I actually made a typo in my process, but yes overall I did it again on paper and it is -6 :)